Two cars collide with each other. The initial speeds are 3 m / s and 2 m / s. Their respective masses are 100kg and 90kg. What will be the car speed of 90kg after the collision.
Answer #1

m1 = mass of first car = 100 kg

v1i = initial velocity of first car = 3 m/s

v1f = final velocity of first car = ?

m2 = mass of second car = 90 kg

v2i = initial velocity of second car = 2 m/s

v2f = final velocity of second car = ?

Using conservation of momentum

m1 v1i + m2 v2i = m1 v1f + m2 v2f

(100) (3) + (90) (2) = (100) v1f + (90) v2f

(100) v1f + (90) v2f = 480

v2f = (48 - 9 v2f )/10                                                                    Eq-1

Using conservation of kinetic energy

m1 v21i + m2 v22i = m1 v21f + m2 v22f

(100) (3)2 + (90) (2)2 = (100) v21f + (90) v22f

Using eq-1

(100) (3)2 + (90) (2)2 = (100) v21f + (90) ((48 - 9 v2f )/10)2

v1f = 1.66 m/s

Using eq-1

v2f = (48 - 9 v2f )/10

v2f = (48 - 9 (1.66))/10

v2f = 3.31 m/s